NvidiaIntely (nvidiaintely):

My Flower Basket delivers flowers within a 25 mile radius of their store. If this area is represented graphically, with My Flower Basket located at (4, 3), what is the equation that represents the delivery area?

OpenStudy (kevin):

(x-4)^2+(y-3)^2=25^2 (x-4)^2+(y-3)^2=625

NvidiaIntely (nvidiaintely):

Thanks!

OpenStudy (kevin):

yw